OpenStudy (anonymous):

rewrite the function in intercept form y=x^2+8x+15

OpenStudy (anonymous):

Intercept form is for lines, and that is the equation for a parabola. Do you need help finding the points where this parabola intercepts the x-axis?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

To do that, you need to set it all to zero and factor: x^2+8x+15 = 0 (x+5)(x+3) = 0 then put both those in parentheses chunks in their own equations x + 5 = 0 x + 3 = 0 and then simplify x = -5 x = -3 So the two points where the parabola crosses the x-axis are -5 and -3.

OpenStudy (anonymous):

Or, since we set y equal to zero in the beginning, in coordinate form it would be (-5,0) and (-3, 0)